# Revisiting Non-Acyclic GFlowNets in Discrete Environments

### v1

@@ Acknowledgements
We would like to thank Leo Maxime Brunswic for the helpful discussion and providing implementation details of the paper (Brunswic et al., 2024). This research was supported in part through computational resources of HPC facilities at HSE University (Kostenetskiy et al., 2021).
@@ Appendix A Proofs
Consider a random walk on $\mathcal{G}$ with reversed edges and backward policy as a transition probability. In particular, we define a stochastic process $\{X_{n}\}_{n=0}^{\infty}$, such that $X_{0}=s_{f}$ a.s. and $\mathbb{P}[X_{i}=s\mid X_{i-1}=s^{\prime}]=\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})$ and also $\mathbb{P}[X_{i}=s_{0}\mid x_{i-1}=s_{0}]=1$, i.e., $s_{0}$ is an absorbing state. We want to show that
@@
    The random walk terminates with probability 1: $\mathbb{P}[\exists n:X_{n}=s_{0}]=1$;
@@
$\sum_{\tau\in\mathcal{T}}\mathbb{P}[(X_{n_{\tau}+1},\ldots,X_{0})=\tau]=\mathbb{P}[\exists\tau\in\mathcal{T}:(X_{n_{\tau}+1},\ldots,X_{0})=\tau]]=\mathbb{P}[\exists n:X_{n}=s_{0}]\,,$
since the family of events $\{(X_{n_{\tau}+1},\ldots,X_{0})=\tau\}$ do not intersect for different $\tau$.
Let us consider any intermediate state $s\in\mathcal{S}\setminus\{s_{0},s_{f}\}$ and the same random walk $\{Y_{n}\}_{n=0}^{\infty}$ that starts from a state $s$. We define $p_{s}\triangleq\mathbb{P}[\exists n>0:Y_{n}=s]$, i.e. the probability that the random walk returns to $s$ at some point. First, notice that $p_{s}<1$. Indeed, based on our assumptions about $\mathcal{G}$, there exists at least one path $\tau$ from $s_{0}$ to $s$, and furthermore, we can only consider paths that do not contain any cycles, as these can be effectively eliminated. In this case, we have
@@
by the condition on $\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})>0$ for $s\to s^{\prime}\in\mathcal{E}$. Notice that $\{(Y_{n_{\tau}+1},\ldots,Y_{0})=\tau\}\cap\{\exists n>0:Y_{n}=s\}=\emptyset$, since if the trajectory of the random walk has already reached $s_{0}$, it will never return. Thus, $p_{s}=\mathbb{P}[\exists n>0:Y_{n}=s]<1$.
Next, for each state $s\in\mathcal{S}\setminus\{s_{0},s_{f}\}$ we define $N_{s}\triangleq\sum_{i=0}^{\infty}\mathbb{I}\{X_{i}=s\}$ as a number of visits of a state $s$ during the original backward random walk, and also let us define $N^{\prime}_{s}\triangleq\sum_{i=0}^{\infty}\{Y_{i}=s\}$ a number of visits of a state $s$ during the backward random walk that starts at $s$. We notice that $\mathbb{E}[N_{s}]\leq\mathbb{E}[N_{s}^{\prime}]$ since after the first visit of $s$, by Markovian property we can replace a random walk $\{X_{n}\}$ with a random walk $\{Y_{n}\}$. At the same time, we have $\mathbb{P}[N^{\prime}_{s}>k]=\mathbb{P}[\exists n_{1},\ldots,n_{k}>0:Y_{n_{j}}=s]$ and, by Markovian property, we have $\mathbb{P}[N^{\prime}_{s}>k]=p_{s}^{k}$. Thus
@@
Firstly, we prove flow matching conditions. We have
@@
$\begin{split}\mathcal{F}(s\to s^{\prime})&=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\mathbb{E}_{\tau}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{I}\{s_{t}=s,s_{t+1}=s^{\prime}\}\mid n_{\tau}}\right]}\right]\\ &=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{E}_{\tau}\mathopen{}\mathclose{{}\left[\mathbb{I}\{s_{t}=s,s_{t+1}=s^{\prime}\}\mid n_{\tau}}\right]}\right]\\ &=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{P}\mathopen{}\mathclose{{}\left(s_{t}=s,s_{t+1}=s^{\prime}\mid n_{\tau}}\right)}\right]\\ &=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{P}\mathopen{}\mathclose{{}\left(s_{t+1}=s^{\prime}\mid n_{\tau}}\right)\mathbb{P}\mathopen{}\mathclose{{}\left(s_{t}=s\mid s_{t+1}=s^{\prime},n_{\tau}}\right)}\right]\\ &=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{P}\mathopen{}\mathclose{{}\left(s_{t+1}=s^{\prime}\mid n_{\tau}}\right)}\right]\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})=\mathcal{F}(s^{\prime})\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime}).\\ \end{split}$
In the penultimate equation we used Markovian property of trajectory distribution.
Finally, ${\mathcal{F}}(s_{0})$ is equal to, by definition, $\mathcal{F}(s_{f})$ multiplied by the expected number of times $\tau\sim\mathcal{P}$ visits $s_{0}$, where the latter is always $1$, so we have ${\mathcal{F}}(s_{0})=\mathcal{F}(s_{f})$.
@@
Since $\mathcal{F}(s\to s^{\prime})$ satisfies flow matching conditions, we will define
@@
Next, take $\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})=\mathcal{F}(s\to s^{\prime})/\mathcal{F}(s^{\prime})$. Let us denote $\hat{\mathcal{F}}$ to be flows from Definition 3.5 that are induced by $\mathcal{P}_{\mathrm{B}}$ and $\mathcal{F}(s_{f})$ (which correspond to expected number of visits in regard to the trajectory distribution $\mathcal{P}$ induced by $\mathcal{P}_{\mathrm{B}}$). In essence, we want to prove that $\mathcal{F}$ and $\hat{\mathcal{F}}$ coincide.
By Proposition 3.6 and definition of $\mathcal{P}_{\mathrm{B}}$ we have
^$\hat{\mathcal{F}}(s\to s^{\prime})=\hat{\mathcal{F}}(s^{\prime})\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})=\frac{\hat{\mathcal{F}}(s^{\prime})}{\mathcal{F}(s^{\prime})}\mathcal{F}(s\to s^{\prime})=C(s^{\prime})\mathcal{F}(s\to s^{\prime}).$
Where we denote $C(s)=\hat{\mathcal{F}}(s)/\mathcal{F}(s)$. In addition, by Proposition 3.6, $\hat{\mathcal{F}}$ satisfies flow matching conditions, thus
@@
The first equation is by definition of $C(s)$ and the second equation is due to flow matching conditions. Thus we have a system of linear equations with respect to $C(s)$:
@@
$\begin{split}\mathcal{F}(s\to s^{\prime})&=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\mathbb{E}_{\tau}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{I}\{s_{t}=s,s_{t+1}=s^{\prime}\}\mid n_{\tau}}\right]}\right]\\ &=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{E}_{\tau}\mathopen{}\mathclose{{}\left[\mathbb{I}\{s_{t}=s,s_{t+1}=s^{\prime}\}\mid n_{\tau}}\right]}\right]\\ &=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{P}\mathopen{}\mathclose{{}\left(s_{t}=s,s_{t+1}=s^{\prime}\mid n_{\tau}}\right)}\right]\\ &=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{P}\mathopen{}\mathclose{{}\left(s_{t}=s\mid n_{\tau}}\right)\mathbb{P}\mathopen{}\mathclose{{}\left(s_{t+1}=s^{\prime}\mid s_{t}=s,n_{\tau}}\right)}\right]\\ &=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{P}\mathopen{}\mathclose{{}\left(s_{t}=s\mid n_{\tau}}\right)}\right]\mathcal{P}_{\mathrm{F}}(s^{\prime}\mid s)=\mathcal{F}(s^{\prime})\mathcal{P}_{\mathrm{F}}(s^{\prime}\mid s).\\ \end{split}$
@@
Thus the existence is proven. Finally, detailed balance conditions follow from the proof of uniqueness.
@@
Similarly, since $\mathcal{P}_{\mathrm{B}}(\cdot\mid s)$ is a valid probability distribution over $\operatorname{in}(s)$, and $\mathcal{F}$, $\mathcal{P}_{\mathrm{F}}$ and $\mathcal{P}_{\mathrm{B}}$ satisfy detailed balance conditions, we have
@@
Thus $\bar{\mathcal{F}}$ satisfies flow matching conditions. In addition
@@
Next, consider any trajectory $\tau=(s_{0},s_{1},\dots,s_{n_{\tau}},s_{f})\in\mathcal{T}$. By detailed balance conditions we have
@@
Then, for any policy $\pi$, we define the regularized value function as follows
$V^{\pi}_{\lambda}(s)\triangleq\mathbb{E}_{\tau\sim\mathsf{P}_{\mathcal{T}}^{\pi}}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{\infty}r(s_{t},s_{t+1})+\lambda\mathcal{H}(\pi(\cdot|s_{t}))\mid s_{0}=s}\right]\,,$ (13)
where $\mathcal{H}$ is Shannon entropy, $\mathsf{P}_{\mathcal{T}}^{\pi}$ is a trajectory distribution induced by the following the policy $\pi$: $s_{t}\sim\pi(\cdot|s_{t-1})$ for all $t\geq 1$ and the starting state $s_{0}$ is fixed as $s$ (not to be confused with the initial state in $\mathcal{G}$). Overall, it is not clear if the value function is a well-defined function in the case of $\gamma=1$. We call this type of problem regularized shortest path.
@@
Assume that (i) a graph $\mathcal{G}$ satisfies 3.1 and (ii) for any $s\in\mathcal{S},s^{\prime}\in\mathcal{A}_{s}$ it holds $r(s,s^{\prime})\leq 0$ and $r(s,s^{\prime})=0$ if and only if $|\mathcal{A}_{s}|=1$. Then, a regularized shortest path problem admits a unique solution, and the value of its solution satisfies soft optimal Bellman equations
@@
Without loss of generality, we can assume that for all states, it holds $|\mathcal{A}_{s}|\geq 2$, because we can shrink edges such that $|\mathcal{A}_{s}|=1$ since they will provide zero reward and zero entropy. Thus, we can assume that $r(s,s^{\prime})<0$ for all $s\in\mathcal{S},s^{\prime}\in\mathcal{A}_{s}$.
Then, we can rewrite the value function in the following way
$V^{\pi}_{\lambda}(s_{0})=\sum_{s\in\mathcal{S}}\sum_{s^{\prime}\in\operatorname{out}(s)}\mathbb{E}_{\pi}[n_{\tau}(s,s^{\prime})]r(s,s^{\prime})+\lambda\sum_{s\in\mathcal{S}}\mathbb{E}_{\pi}[n_{\tau}(s)]\mathcal{H}(\pi(\cdot|s))\,.$
where $n_{\tau}(s)=\sum_{t=0}^{\infty}\mathbb{I}\{s_{t}=s\}$ and $n_{\tau}(s,s^{\prime})=\sum_{t=0}^{\infty}\mathbb{I}\{s_{t}=s,s_{t+1}=s^{\prime}\}$ is a number of visits of a particular state and edge in $\mathcal{G}$. In the analogy with occupancy measures in RL, we employ the notation $d^{\pi}(s)=\mathbb{E}_{\pi}[n_{\tau}(s)]$ and $d^{\pi}(s,s^{\prime})=\mathbb{E}_{\pi}[n_{\tau}(s,s^{\prime})]$ for an expected number of visits. However, this definition also corresponds to the flow function in the reversed graph (see Definition 3.5) with the ”backward policy” $\pi$.
Given this expression, we notice that if we consider any policy such that $\pi(s^{\prime}|s)>0$ for all $s\in\mathcal{S}$, then, according to Lemma 3.4 in the reversed graph, $d^{\pi}(s,s^{\prime}),d^{\pi}(s)<+\infty$, and it implies that the value function is finite.
Additionally, any policy such that $\mathbb{E}_{\pi}[n_{\tau}(s,s^{\prime})]=+\infty$ have $V^{\pi}_{\lambda}(s_{0})=-\infty$ and thus cannot be optimal since there exists at least one policy with a finite value. Therefore, we can consider only a class of policies such that $d^{\pi}(s,s^{\prime})\in[0,+\infty)$.
Then, we notice that $d^{\pi}(s,s^{\prime})=\pi(s^{\prime}|s)d^{\pi}(s)$ thus we can rewrite the value in the following form
$V^{\pi}_{\lambda}(s_{0})=\sum_{s\in\mathcal{S}}\sum_{s^{\prime}\in\operatorname{out}(s)}d^{\pi}(s,s^{\prime})r(s,s^{\prime})-\underbrace{\lambda\sum_{s\in\mathcal{S}}\sum_{s^{\prime}\in\operatorname{out}(s)}d^{\pi}(s,s^{\prime})\log\mathopen{}\mathclose{{}\left(d^{\pi}(s,s^{\prime})/\sum_{s^{\prime}\in\operatorname{out}(s)}d^{\pi}(s,s^{\prime})}\right)}_{R(d^{\pi})}\,.$
As a function of $d^{\pi}(s,s^{\prime})$, we see that the first term in the expression above is linear whereas the second one is relative conditional entropy (Neu et al., 2017) that is strongly concave. Given that the set of all admissible $d^{\pi}(s,s^{\prime})$ is a polytope that is defined as a family of negative functions that satisfies the flow matching constraints (see Proposition 3.6)
@@
Next, we want to prove that the value of the optimal policy satisfies soft optimal Bellman equations. First, we notice that the usual Bellman equations still hold
$Q^{\pi}_{\lambda}(s,s^{\prime})=r(s,s^{\prime})+V^{\pi}_{\lambda}(s^{\prime}),\qquad V^{\pi}_{\lambda}(s)=\sum_{s^{\prime}\in\operatorname{out}(s)}\pi(s^{\prime}|s)Q^{\pi}_{\lambda}(s,s^{\prime})+\lambda\mathcal{H}(\pi(\cdot|s))\,.$
Let us consider a regularized policy improvement operation, defined as
@@
Since the policy $\pi$ is positive, then Lemma 3.4 in the reversed graph implies that $d^{\pi}(s,s^{\prime}),d^{\pi}(s)<+\infty$ and thus all values and Q-values are finite: $Q^{\pi}(s,s^{\prime})>-\infty$ for any $s\in\mathcal{S},s^{\prime}\in\operatorname{out}(s)$ It implies that $\pi^{\prime}$ is also positive. Thus, its trajectories are finite with probability 1 and yields $V^{\pi^{\prime}}_{\lambda}(s_{0})\geq V^{\pi}_{\lambda}(s_{0})$. Finally, applying policy improvement to $\pi^{\star}$ we conclude the statement. ∎

### v2

@@ Acknowledgements
We would like to thank Leo Maxime Brunswic for the helpful discussion and providing implementation details of the paper (Brunswic et al., 2024). This work was supported by the Ministry of Economic Development of the Russian Federation (code 25-139-66879-1-0003). This research was supported in part through computational resources of HPC facilities at HSE University (Kostenetskiy et al., 2021).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Proofs
Consider a random walk on $\mathcal{G}$ with reversed edges transition probabilities given by backward policy. Specifically, we define a Markov chain $\{X_{n}\}_{n=0}^{\infty}$, such that $X_{0}=s_{f}$ a.s. and $\mathbb{P}[X_{i}=s\mid X_{i-1}=s^{\prime}]=\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})$. Let also $\mathbb{P}[X_{i}=s_{0}\mid X_{i-1}=s_{0}]=1$, i.e., $s_{0}$ is an absorbing state. We want to show that
@@
    The random walk terminates at $s_{0}$ with probability 1: $\mathbb{P}[\exists n:X_{n}=s_{0}]=1$;
@@
$\sum_{\tau\in\mathcal{T}}\mathbb{P}[(X_{n_{\tau}+1},\ldots,X_{0})=\tau]=\mathbb{P}[\exists\tau\in\mathcal{T}:(X_{n_{\tau}+1},\ldots,X_{0})=\tau]=\mathbb{P}[\exists n:X_{n}=s_{0}]\,,$
since the events $\{(X_{n_{\tau}+1},\ldots,X_{0})=\tau\}$ do not intersect for different $\tau$.
First consider any intermediate state $s\in\mathcal{S}\setminus\{s_{0},s_{f}\}$ and define a Markov chain $\{Y_{n}\}_{n=0}^{\infty}$ with the same transition probabilities as $\{X_{n}\}_{n=0}^{\infty}$, with $Y_{0}=s$ a.s. We define $p_{s}\triangleq\mathbb{P}[\exists n>0:Y_{n}=s]$, i.e. the probability that $\{Y_{n}\}$ returns to $s$. First, notice that $p_{s}<1$. Indeed, based on our assumptions on $\mathcal{G}$, there exists at least one path $\tau$ from $s_{0}$ to $s$, and furthermore, there exists such a path without cycles. In this case, we have
@@
by the condition on $\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})>0$ for $s\to s^{\prime}\in\mathcal{E}$. Notice that $\{(Y_{n_{\tau}+1},\ldots,Y_{0})=\tau\}\cap\{\exists n>0:Y_{n}=s\}=\emptyset$, since if the trajectory of the random walk has already reached $s_{0}$, it will never return to $s$. Thus, $p_{s}=\mathbb{P}[\exists n>0:Y_{n}=s]<1$.
Next, for each state $s\in\mathcal{S}\setminus\{s_{0},s_{f}\}$ we define $N_{s}\triangleq\sum_{i=0}^{\infty}\mathbb{I}\{X_{i}=s\}$ as the number of visits of a state $s$ encountered by the original process. Also, let us define $N^{\prime}_{s}\triangleq\sum_{i=0}^{\infty}\mathbb{I}\{Y_{i}=s\}$ as the number of visits of a state $s$ during the backward random walk that starts at $s$. We notice that
$\mathbb{E}[N_{s}]=\mathbb{P}(\exists k>0:X_{k}=s)\mathbb{E}[N_{s}^{\prime}]\leq\mathbb{E}[N_{s}^{\prime}]\,,$
where the first equation is due to the strong Markov property. At the same time, we have $\mathbb{P}[N^{\prime}_{s}>k]=\mathbb{P}[\exists n_{1},\ldots,n_{k}>0:Y_{n_{j}}=s]$ and, by Markov property, we have $\mathbb{P}[N^{\prime}_{s}>k]=p_{s}^{k}$. Thus
@@
The first statement of the Lemma directly follows from the finiteness of the expected length of the walk, because otherwise it has an infinite length with non-zero probability, leading to a contradiction.
@@
First, we prove the flow matching conditions. We have
@@
$\begin{split}\mathcal{F}(s\to s^{\prime})&=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\mathbb{E}_{\tau}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{I}\{s_{t}=s,s_{t+1}=s^{\prime}\}\mid n_{\tau}}\right]}\right]\\ &=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{E}_{\tau}\mathopen{}\mathclose{{}\left[\mathbb{I}\{s_{t}=s,s_{t+1}=s^{\prime}\}\mid n_{\tau}}\right]}\right]\\ &=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{P}\mathopen{}\mathclose{{}\left(s_{t}=s,s_{t+1}=s^{\prime}\mid n_{\tau}}\right)}\right]\\ &\overset{(a)}{=}\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{P}\mathopen{}\mathclose{{}\left(s_{t+1}=s^{\prime}\mid n_{\tau}}\right)\mathbb{P}\mathopen{}\mathclose{{}\left(s_{t}=s\mid s_{t+1}=s^{\prime},n_{\tau}}\right)}\right]\\ &=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{P}\mathopen{}\mathclose{{}\left(s_{t+1}=s^{\prime}\mid n_{\tau}}\right)}\right]\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})=\mathcal{F}(s^{\prime})\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime}).\\ \end{split}$
Here in (a) we used the Markov property of trajectory distribution. Finally, by definition, ${\mathcal{F}}(s_{0})$ is equal to, $\mathcal{F}(s_{f})$ multiplied by the expected number of times $\tau\sim\mathcal{P}$ visits $s_{0}$, where the latter is always $1$, so we have ${\mathcal{F}}(s_{0})=\mathcal{F}(s_{f})$.
@@
Since $\mathcal{F}(s\to s^{\prime})$ satisfies the flow matching conditions, we define
@@
Next, take $\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})=\mathcal{F}(s\to s^{\prime})/\mathcal{F}(s^{\prime})$. Let us denote $\hat{\mathcal{F}}$ to be flows from Definition 3.5 that are induced by $\mathcal{P}_{\mathrm{B}}$ and $\mathcal{F}(s_{f})$ (which correspond to expected number of visits with respect to the trajectory distribution $\mathcal{P}$ induced by $\mathcal{P}_{\mathrm{B}}$). We aim to prove that $\mathcal{F}$ and $\hat{\mathcal{F}}$ coincide.
By Proposition 3.6 and definition of $\mathcal{P}_{\mathrm{B}}$, we have
^$\hat{\mathcal{F}}(s\to s^{\prime})=\hat{\mathcal{F}}(s^{\prime})\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})=\frac{\hat{\mathcal{F}}(s^{\prime})}{\mathcal{F}(s^{\prime})}\mathcal{F}(s\to s^{\prime})=C(s^{\prime})\mathcal{F}(s\to s^{\prime})\,,$
where we denote $C(s)=\hat{\mathcal{F}}(s)/\mathcal{F}(s)$. In addition, by Proposition 3.6, $\hat{\mathcal{F}}$ satisfies the flow matching conditions, thus
@@
The first equation is by definition of $C(s)$ and the second equation is due to the flow matching conditions. Thus we have a system of linear equations with respect to $C(s)$:
@@
$\begin{split}\mathcal{F}(s\to s^{\prime})&=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\mathbb{E}_{\tau}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{I}\{s_{t}=s,s_{t+1}=s^{\prime}\}\mid n_{\tau}}\right]}\right]\\ &=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{E}_{\tau}\mathopen{}\mathclose{{}\left[\mathbb{I}\{s_{t}=s,s_{t+1}=s^{\prime}\}\mid n_{\tau}}\right]}\right]\\ &=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{P}\mathopen{}\mathclose{{}\left(s_{t}=s,s_{t+1}=s^{\prime}\mid n_{\tau}}\right)}\right]\\ &=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{P}\mathopen{}\mathclose{{}\left(s_{t}=s\mid n_{\tau}}\right)\mathbb{P}\mathopen{}\mathclose{{}\left(s_{t+1}=s^{\prime}\mid s_{t}=s,n_{\tau}}\right)}\right]\\ &=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{P}\mathopen{}\mathclose{{}\left(s_{t}=s\mid n_{\tau}}\right)}\right]\mathcal{P}_{\mathrm{F}}(s^{\prime}\mid s)=\mathcal{F}(s)\mathcal{P}_{\mathrm{F}}(s^{\prime}\mid s).\\ \end{split}$
@@
Thus the existence is proven. Finally, the detailed balance conditions follow from the proof of uniqueness.
@@
Similarly, since $\mathcal{P}_{\mathrm{B}}(\cdot\mid s)$ is a valid probability distribution over $\operatorname{in}(s)$, and $\mathcal{F}$, $\mathcal{P}_{\mathrm{F}}$ and $\mathcal{P}_{\mathrm{B}}$ satisfy the detailed balance conditions, we have
@@
Thus $\bar{\mathcal{F}}$ satisfies the flow matching conditions. In addition
@@
Next, consider any trajectory $\tau=(s_{0},s_{1},\dots,s_{n_{\tau}},s_{f})\in\mathcal{T}$. By the detailed balance conditions we have
@@
Then, for any policy $\pi$, we define the regularized value function for all $s\not=s_{f}$ as follows
$V^{\pi}_{\lambda}(s)\triangleq\mathbb{E}_{\tau\sim\mathsf{P}_{\mathcal{T}}^{\pi}}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}r(s_{t},s_{t+1})+\lambda\mathcal{H}(\pi(\cdot|s_{t}))\mid s_{0}=s}\right]\,,$ (13)
and $V^{\pi}_{\lambda}(s_{f})=0$, where $\mathcal{H}$ is Shannon entropy, $\mathsf{P}_{\mathcal{T}}^{\pi}$ is a trajectory distribution induced by the following the policy $\pi$: $s_{t}\sim\pi(\cdot|s_{t-1})$ for all $t\geq 1$ and the starting state $s_{0}$ is fixed as $s$ (not to be confused with the initial state in $\mathcal{G}$), and $n_{\tau}$ is a length of trajectory defined as $n_{\tau}=\min\{k\geq 0\mid s_{k+1}=s_{f}\}$. Overall, it is not clear if the value function is a well-defined function when no discounting is used ($\gamma=1$). We call this problem a regularized shortest path problem, akin to shortest path and stochastic shortest path problem (Bertsekas, 2012, Chapter 3). A policy $\pi^{\star}$ is called optimal if it maximizes $V^{\pi}_{\lambda}(s_{0})$.
@@
Assume that (i) a graph $\mathcal{G}$ satisfies 3.1 and (ii) for any $s\in\mathcal{S},s^{\prime}\in\mathcal{A}_{s}$ it holds $r(s,s^{\prime})\leq 0$ and $r(s,s^{\prime})=0$ if and only if $|\operatorname{in}(s^{\prime})|=1$. Also, assume that for any optimal policy $\pi^{\star}$ it holds $\mathbb{E}_{\pi^{\star}}[n_{\tau}]<+\infty$.
Then, a regularized shortest path problem admits a unique solution, and the value of its solution satisfies soft optimal Bellman equations
@@
Let us define a number of visits of a vertex $s$ and an edge $s,s^{\prime}$ in $\mathcal{G}$ on a given trajectory $\tau$ as $n_{\tau}(s)=\sum_{t=0}^{\infty}\mathbb{I}\{s_{t}=s\}$ and $n_{\tau}(s,s^{\prime})=\sum_{t=0}^{\infty}\mathbb{I}\{s_{t}=s,s_{t+1}=s^{\prime}\}$. In the analogy with occupancy measures in RL, we employ the notation $d^{\pi}(s)\triangleq\mathbb{E}_{\pi}[n_{\tau}(s)]$ and $d^{\pi}(s,s^{\prime})\triangleq\mathbb{E}_{\pi}[n_{\tau}(s,s^{\prime})]$ for an expected number of visits. This definition also corresponds to the flow function in the reversed graph (see Definition 3.5) with the ”backward policy” $\pi$. The condition on expected trajectory length of optimal policies implies that we can consider only policies $\pi$ such that $d^{\pi}(s)<+\infty$ for any $s\in\mathcal{S}\setminus\{s_{0},s_{f}\}$, and, as a result, $d^{\pi}(s,s^{\prime})<+\infty$.
Next, we rewrite the value function in the initial state as follows
$V^{\pi}_{\lambda}(s_{0})=\sum_{s\in\mathcal{S}\setminus\{s_{f}\}}\sum_{s^{\prime}\in\operatorname{out}(s)}d^{\pi}(s,s^{\prime})r(s,s^{\prime})+\lambda\sum_{s\in\mathcal{S}}d^{\pi}(s)\mathcal{H}(\pi(\cdot|s))\,.$
Then, we notice that $d^{\pi}(s,s^{\prime})=\pi(s^{\prime}|s)\cdot d^{\pi}(s)$ thus we can rewrite the value in the following form
$V^{\pi}_{\lambda}(s_{0})=\sum_{s\in\mathcal{S}\setminus\{s_{f}\}}\sum_{s^{\prime}\in\operatorname{out}(s)}d^{\pi}(s,s^{\prime})r(s,s^{\prime})-\underbrace{\lambda\sum_{s\in\mathcal{S}}\sum_{s^{\prime}\in\operatorname{out}(s)}d^{\pi}(s,s^{\prime})\log\mathopen{}\mathclose{{}\left(d^{\pi}(s,s^{\prime})/\sum_{s^{\prime}\in\operatorname{out}(s)}d^{\pi}(s,s^{\prime})}\right)}_{R(d^{\pi})}\,.$
As a function of $d^{\pi}(s,s^{\prime})$, we see that the first term in the expression above is linear whereas the second one is relative conditional entropy (Neu et al., 2017) that is strongly concave. Given that the set of all admissible $d^{\pi}(s,s^{\prime})$ is a polytope that is defined as a family of negative functions that satisfies the flow matching conditions (see Proposition 3.6)
@@
Next, we want to prove that the value of the optimal policy satisfies soft optimal Bellman equations. First, we notice that the usual Bellman equations still hold since $n_{\tau}$ is a stopping time
$Q^{\pi}_{\lambda}(s,s^{\prime})=r(s,s^{\prime})+V^{\pi}_{\lambda}(s^{\prime}),\qquad V^{\pi}_{\lambda}(s)=\sum_{s^{\prime}\in\operatorname{out}(s)}\pi(s^{\prime}|s)Q^{\pi}_{\lambda}(s,s^{\prime})+\lambda\mathcal{H}(\pi(\cdot|s))\,,$
with an additional initial condition $V^{\pi}_{\lambda}(s_{f})=0$, by the proprieties of conditional expectation. Let us consider a regularized policy improvement operation, defined as
@@
Since the policy $\pi$ is positive, then Lemma 3.4 in the reversed graph implies that $d^{\pi}(s,s^{\prime}),d^{\pi}(s)<+\infty$ and thus all values and Q-values are finite: $Q^{\pi}(s,s^{\prime})>-\infty$ for any $s\in\mathcal{S},s^{\prime}\in\operatorname{out}(s)$. It implies that $\pi^{\prime}$ is also positive. Thus, its trajectories are finite with probability 1 and yields $V^{\pi^{\prime}}_{\lambda}(s_{0})\geq V^{\pi}_{\lambda}(s_{0})$. Finally, applying policy improvement to $\pi^{\star}$ we conclude the statement. ∎